OpenStudy (anonymous):

Find the linear function f(x) such that f(2)=-3 and f(6)=-5.

OpenStudy (anonymous):

f(x)=0.5x-2

OpenStudy (anonymous):

that doesn't work.

OpenStudy (anonymous):

f(x)=0.5x-2 works...

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

no problem. nothing to be sorry about! :)

OpenStudy (anonymous):

I worked it by finding the slope, and then substituting the values given in y = mx+c

OpenStudy (anonymous):

Hes right, it doesnt work.

OpenStudy (anonymous):

Would any of you mind showing me how you got there? Thanks :-)

OpenStudy (anonymous):

the correct answer is f(x)=-1/2x-2 but I don't know how to get there.

OpenStudy (anonymous):

go with what mandar suggested

OpenStudy (anonymous):

elizated forgot to add the negative sign in his equation

OpenStudy (anonymous):

Since it's a linear equation, the equation is a straight line. The points given are (2, -3) and (6, -5) Find the slope, and use y = mx+c -> f(x) = -0.5x +c Find c using the values given in the problem statement.

OpenStudy (anonymous):

well set up an x y chart: x | y 0 - 1 - 2 -3 3 - 4 - 5 - 6 -5 then start filling in the missing numbers probably starting with 4: -4. since it's counting by 0.5, that's your slope. now wor back to zero and that's your y intercept.

OpenStudy (anonymous):

woops, sorry it is negative. and i'm a girl by the way...

OpenStudy (anonymous):

Thank you very much!